What is the equation of the line that is parallel to 3x+y=-18 that passes through (46, -49)? Please I really need help!
Answer by jim_thompson5910(35256)

Anything parallel to 3x+y=-18 is of the form 3x+y=C where C is any number. To find the desired equation, we simply plug in x = 46 and y = -49 to find the correct value of C to find the answer.


3x+y=C


3(46)+(-49)=C


138-49=C


89=C


C = 89


So the answer is 3x+y = 89